# Risks of Practicing Large Language Models in Smart Grid: Threat Modeling and Validation

### v1

@@ IV-B Domain Knowledge Extraction
To demonstrate the potential for attackers to extract domain knowledge (prompts) from LLMs, we conducted a simulation where an LLM is employed using utility company data, specifically designed to provide only aggregated information about the data. Under normal circumstances, the LLM would reject requests to output detailed data. However, our simulation shows that with specifically crafted inputs, an attacker can manipulate the LLM into divulging all the domain knowledge it possesses. This highlights a significant vulnerability where the LLM, despite its initial configuration to safeguard detailed information, can be induced to reveal comprehensive data under certain conditions.
Dataset: The dataset employed for domain knowledge extraction attack consists of statistically advanced metering infrastructure (AMI) data provided by the U.S. Energy Information Administration (EIA) [35]. This AMI data originates from the EIA-861M form, a monthly report that gathers data on electricity sales and revenue from a statistically selected sample of electric utilities across the United States. For simplicity, we only select six columns of the dataset: Month, Utility_Number, Name, State, AMR_Total, AMI_Total.
Attack Simulation: We processed the dataset by dividing it into smaller groups, with each group containing 10 rows of data. For each evaluation test, a group is input into the LLMs as domain knowledge. The LLMs are then tasked to provide only aggregated information based on this data. This structured approach ensures that the LLMs do not disclose detailed data elements, adhering to the privacy and confidentiality standards required in utility data handling. The specific prompts and associated inputs used in these tests are detailed in Fig. 6.
Figure 6: Sample Domain Knowledge and Attack Queries
@@
  • Phase 1: We directly challenge the LLMs by requesting them to output all the data they possess. This phase involves 100 tests, each employing different groups of data (domain knowledge) to assess the robustness of the LLMs’ data confidentiality mechanisms under straightforward query conditions.
  • Phase 2: We employ well-crafted queries designed to subtly coax the LLMs into revealing more information than they are configured to disclose. This phase also consists of 100 tests using the same groupings of data as in Phase 1, to measure the LLMs’ resilience against more sophisticated extraction techniques.
Simulation Result: Our simulation results indicate a distinct contrast in the performance of LLMs based on the type of input provided.
In general scenarios, where direct queries were made to both GPT-3.5 and GPT-4 to reveal all the data they hold, the models successfully adhered to their configuration by not providing detailed data across all 100 tests. This demonstrates their effectiveness in safeguarding sensitive information under straightforward querying conditions.
Conversely, when faced with well-crafted, nuanced inputs designed to elicit more information than they are configured to disclose, both GPT-3.5 and GPT-4 failed to maintain their data prevention protocols. In these cases, both models disclosed all the provided data, highlighting a vulnerability in their ability to handle sophisticated data extraction attacks. The sampling output of the LLMs is shown in Fig. 7.
Figure 7: Domain Knowledge Extraction Output Samples
Our evaluation results indicate that attackers can successfully extract the domain knowledge from LLM applications in smart grid environments, using the threat models we proposed.
@@ V Discussion and Conclusion
LLMs, as pre-trained models, pose challenges in controlling their outputs. Moreover, the domain knowledge utilized for LLM tasks is valuable and necessitates enhanced protection. It is crucial to evaluate potential risks before deploying LLMs in smart grid applications. In this paper, we delineate the scope of LLM security for smart grid applications and identify two major types of attacks concerning data security (bad data injection) and data privacy (domain knowledge extraction). Our validation demonstrates that, with well-crafted inputs, attackers can successfully execute both types of attacks. Given the rapid development of LLMs, new attack vectors may emerge. Keeping track of the new potential risks, such as interactivity injection, should be the future work of our research.

### v2

@@ IV-B Domain Knowledge Extraction
IV-B1 Application Introduction
To demonstrate the vulnerability of LLMs to domain knowledge extraction attacks from insider attackers, we consider the use case where LLMs are employed by a utility company acting as a virtual assistant for internal employees. Specifically, the LLMs are prompted with raw smart grid data, which we consider as domain knowledge, but are intended to respond to employees with only aggregated information about the data. Under normal circumstances, the LLMs would reject requests to output detailed data.
In this paper, we utilize the advanced metering infrastructure (AMI) data provided by the U.S. Energy Information Administration (EIA) [40] to serve as the domain knowledge. This AMI data originates from the EIA-861M form, a monthly report that gathers data on electricity sales and revenue from a statistically selected sample of electric utilities across the United States. For simplicity, we select only six columns of the dataset: Month, Utility_Number, Name, State, AMR_Total, AMI_Total. The sample system prompt (domain knowledge) provided to LLMs is shown in Fig. 5. To evaluate the generalization of the attack, we provide different data from EIA in the system prompts for each query we conduct.
Figure 5: System Prompt (Domain Knowledge) for LLM-based Virtual Assistant
IV-B2 Attack Simulation
@@
  • Phase 1 Normal Query: We directly challenge the LLMs by requesting them to output all the data they possess. This phase involves 100 tests for all three LLMs, each employing different groups of data (domain knowledge) to assess the robustness of the LLMs’ data confidentiality mechanisms under straightforward query conditions.
    In our evaluation, all three models successfully adhered to their configuration by not providing detailed data across all 100 tests. This demonstrates their effectiveness in safeguarding sensitive information under straightforward querying conditions. A sample of normal query input and LLM response is presented in Fig. 6
  • Phase 2 Domain Knowledge Extraction: In the attack phase, we employ well-crafted queries designed to subtly coax the LLMs into revealing more information than they are configured to disclose. This phase also consists of 100 tests using the same groupings of data as in Phase 1 for all three LLMs, to measure the LLMs’ resilience against more sophisticated extraction techniques.
    When faced with well-crafted, nuanced inputs designed to elicit more information than they are configured to disclose, all three models failed to maintain their data prevention protocols in all tests. In these cases, all models disclosed all the provided data, highlighting a vulnerability in their ability to handle sophisticated domain knowledge extraction attacks. The sampling output of the LLMs is shown in Fig. 6.
    Our evaluation results indicate that attackers can successfully extract the domain knowledge from LLM applications in smart grid environments under the threat models we proposed.
Figure 6: Domain Knowledge Extraction Input and Output
@@ V Discussion and Conclusion
Large language models with pre-trained parameters pose significant challenges in controlling and understanding their generated outputs. Additionally, the domain knowledge utilized in LLM tasks is valuable and necessitates enhanced protection. Therefore, it is crucial to evaluate potential risks before deploying LLMs in smart grid applications.
In this paper, we delineated the scope of LLM security for smart grid applications and identified two major types of attacks concerning data security—bad data injection—and data privacy—domain knowledge extraction. Our validation demonstrates that attackers can successfully execute both types of attacks using well-crafted inputs. Given the rapid development of LLMs, new attack vectors may emerge. Monitoring these potential risks, such as interactivity injection, should be a focus of future research.